OpenStudy (ksaimouli):

a compact disc has a radius of 6cm if the disc rotates about its central axis at a constant angular speed of 5 rev/s what is the total distance traveled by a point on the rim of the disc 40min

OpenStudy (ksaimouli):

when i did \[\omega=\frac{ \theta }{ t }\]

OpenStudy (ksaimouli):

but why? (ur answer)

OpenStudy (ksaimouli):

it did not work

OpenStudy (shane_b):

This question doesn't even require physics...it's purely mathematical: First, calculate the circumference of the circle. Then multiply that by 5 since it makes 5 revolutions per second. That will give you how far the dot traveled in 1 second. Then just multiply that times how many seconds are in 40 minutes.

OpenStudy (ksaimouli):

Can u show me in physics way

OpenStudy (vincent-lyon.fr):

\(\omega=\frac{ \theta }{ t }\) is ok. You should just be careful to plug in the values in the right units.

OpenStudy (ksaimouli):

but @Vincent-Lyon.Fr i did not get the answer with this equation

OpenStudy (vincent-lyon.fr):

Show us you numbers.

OpenStudy (ksaimouli):

\[\theta=10\pi(2400)\]

OpenStudy (vincent-lyon.fr):

What the question asks is the DISTANCE travelled, not the angle swept during that time.

OpenStudy (ksaimouli):

Distance in rotation is nothing but the angle right?

OpenStudy (vincent-lyon.fr):

No, distance is distance (in metres), and angle is angle (in radians).

OpenStudy (ksaimouli):

so i found theta so can i use s=r theta

OpenStudy (ksaimouli):

\[s=r \theta\]

OpenStudy (vincent-lyon.fr):

Yep and you should find about 4.5 km